### v1

@@ 0.5.3 Verification Pipeline Setup and the Role of Embedding Functions
  • semantic subspaces should be preferred over geometric subpspaces;
  • best semantic subspaces result from using stronger NLP perturbations;
  • adversarial training should be done on these “best” subpspaces;
  • Marabou performs best thanks to its precision.
@@
Based on this, we further strengthen NLP perturbations: we substitute Polyjuice used in the previous section with Vicuna. Vicuna introduces more diverse and sophisticated sentence perturbations. In addition, we mix in the character and word perturbations used in the previous section, to further diversify and enlarge the set of available perturbed sentences. In the terminology of Section 0.4.1, we obtain the sets of perturbed sentences $\mathcal{A}^{pos}_{pert}$ and $\mathcal{A}^{neg}_{pert}$ where $pert$ is itself a combined kind of perturbations, i.e. $pert=\{char,word,vicuna\}$. Table 18 also uses notation $\mathcal{A}^{pos}_{pert^{\diamondsuit}}$ and $\mathcal{A}^{neg}_{pert^{\diamondsuit}}$ to refer to filtered sets, this terminology will be introduced in the next subsection. Having $\mathcal{A}^{pos}_{pert}$, the set of hyper-rectangles $\mathbb{H}^{*}_{pert}$ is obtained according to Definition 3. Accordingly, we set the adversarial training to explore hyper-rectangles in $\mathbb{H}^{*}_{pert}$, and obtain the adversarially robust DNN $N_{pert}$.
@@
Pipeline Component Section 0.5 Choices Additional Details
0. Choice of Datasets
RUAR, Medical
Same as in Section 0.4 experiments. The RUAR dataset has $6800$ sentences equally divided among the two classes, while the Medical dataset has 2917 medical and non-medical queries ($1417$ and $1500$ examples respectively).
1. Generating Sentence Perturbations
$\mathcal{A}^{pos}_{pert}$, $\mathcal{A}^{pos}_{pert^{\diamondsuit}}$, $\mathcal{A}^{neg}_{pert}$, $\mathcal{A}^{neg}_{pert^{\diamondsuit}}$
With $pert=\{char,word,vicuna\}$, the resulting set of sentences $\mathcal{A}_{pert}$ has $54400$ sentences for RUAR and $15824$ sentences for Medical. The superscript ${}^{\diamondsuit}$ refers to filtering that will be introduced in Section 0.5.4.
2. Embedding Sentences into Real Vector Space
s-bert 22M, s-gpt 1.3B, s-gpt 2.7B
In the experiments of Section 0.4 only s-bert 22M was used.
3. Defining Semantic Subspaces based on Sentence Perturbations
$\mathbb{H}^{*}_{pert}$, $\mathbb{H}^{*}_{pert^{\diamondsuit}}$
$\mathbb{H}^{*}_{pert}$ and $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ are obtained on $\mathcal{A}^{pos}_{pert}$ and $\mathcal{A}^{pos}_{pert^{\diamondsuit}}$, respectively. Their cardinality is $3400$ for RUAR and $989$ for Medical.
  • Volume of $\mathbb{H}^{*}_{pert}$ for RUAR is $191.83e-$ 19 (s-bert 22M), $353.24e+$ 35 (s-gpt 1.3B), $363.30e+$ 36 (s-gpt 2.7B).
  • Volume of $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ for RUAR is $252.43e-$ 25 (s-bert 22M), $271.54e+$ 27 (s-gpt 1.3B), $283.10e+$ 28 (s-gpt 2.7B).
  • Volume of $\mathbb{H}^{*}_{pert}$ for Medical is $223.13e-$ 22 (s-bert 22M), $331.70e+$ 33 (s-gpt 1.3B), $332.10e+$ 33 (s-gpt 2.7B).
  • Volume of $\mathbb{H}^{*}_{pert^{\diamondsuit}}$ for Medical is $283.65e-$ 28 (s-bert 22M), $253.30e+$ 25 (s-gpt 1.3B), $273.83e+$ 27 (s-gpt 2.7B).
4. Training Robust DNNs using Semantic Subspaces
$N_{base}$, $N_{pert}$, $N_{pert^{\diamondsuit}}$
$N_{base}$ is obtained as in Section 0.4, while $N_{pert}$ and $N_{pert^{\diamondsuit}}$ are obtained through our adversarial training on $\mathbb{H}^{*}_{pert}$ and $\mathbb{H}^{*}_{pert{\diamondsuit}}$, respectively.
5. Verifying resulting DNNs on the given semantic subpspaces
Marabou
Same settings as in Section 0.4
Table 18: Section 0.5 NLP verification pipeline setup, implemented using ANTONIO. Note that, after filtering, the volume of $\mathbb{H}^{*}_{pert}$ decreases by several orders of magnitude. Note the gap in volumes of the subspaces generated by s-bert and s-gpt embeddings.
For illustration as well as the initial sanity check, we report accuracy of the obtained models, per each of the chosen embedding functions, in Table 19. Overall the figures are as expected: the choice of the encoder influences accuracy. Thus, using an additional DNN as a filter (as opposed to the zero-shot model), dramatically increases the accuracy of the overall system, from $54-64\%$ in the previous section to the range of $76-95\%$ in Table 19.
@@ 5. Verifying resulting DNNs on the given semantic subpspaces
RUAR $N_{zero-shot}$ 0.5167 0.5835 0.5481 - - -
$N_{base}$ (s-bert 22M) 0.9568 0.9129 0.9344 0.9477 0.7186 0.8174
$N_{pert}$ (s-bert 22M) 0.8497 0.9863 0.9129 0.8125 0.9466 0.8745
$N_{base}$ (s-gpt 1.3B) 0.9620 0.8725 0.9151 0.9545 0.6738 0.7898
$N_{pert}$ (s-gpt 1.3B) 0.6303 0.9980 0.7724 0.6126 0.9860 0.7554
$N_{base}$ (s-gpt 2.7B) 0.9674 0.8729 0.9177 0.9549 0.6982 0.8066
$N_{pert}$ (s-gpt 2.7B) 0.6018 0.9980 0.7508 0.5846 0.9899 0.7350
Medical $N_{zero-shot}$ 0.5895 0.7022 0.6409 - - -
$N_{base}$ (s-bert 22M) 0.9523 0.9325 0.9423 0.9520 0.8964 0.9234
$N_{pert}$ (s-bert 22M) 0.9335 0.9736 0.9531 0.9238 0.9517 0.9376
$N_{base}$ (s-gpt 1.3B) 0.9193 0.8811 0.8998 0.9217 0.8417 0.8798
$N_{pert}$ (s-gpt 1.3B) 0.8441 0.9627 0.8938 0.8315 0.9470 0.8854
$N_{base}$ (s-gpt 2.7B) 0.9325 0.8929 0.9123 0.9289 0.8479 0.8866
$N_{pert}$ (s-gpt 2.7B) 0.8603 0.9656 0.9098 0.8488 0.9499 0.8964
Table 19: Accuracy of the models on the test/perturbation set. The average standard deviation is $0.0049$.
Dataset Model Test set Perturbed test set
Precision Recall F1 Precision Recall F1
RUAR $N_{base^{\text{ \ }}}$ (s-bert 22M) 0.9568 0.9129 0.9344 0.9477 0.7186 0.8174
$N_{pert^{\diamondsuit}}$ (s-bert 22M) 0.8507 0.9894 0.9148 0.8289 0.9412 0.8815
$N_{base^{\text{ \ }}}$ (s-gpt 1.3B) 0.9620 0.8725 0.9151 0.9545 0.6738 0.7898
$N_{pert^{\diamondsuit}}$ (s-gpt 1.3B) 0.6493 0.9965 0.7862 0.6356 0.9808 0.7713
$N_{base^{\text{ \ }}}$ (s-gpt 2.7B) 0.9674 0.8729 0.9177 0.9549 0.6982 0.8066
$N_{pert^{\diamondsuit}}$ (s-gpt 2.7B) 0.6305 0.9969 0.7724 0.6143 0.9852 0.7566
Medical $N_{base^{\text{ \ }}}$ (s-bert 22M) 0.9523 0.9325 0.9423 0.9520 0.8964 0.9234
$N_{pert^{\diamondsuit}}$ (s-bert 22M) 0.9313 0.9717 0.9511 0.9251 0.9493 0.9370
$N_{base^{\text{ \ }}}$ (s-gpt 1.3B) 0.9193 0.8811 0.8998 0.9217 0.8417 0.8798
$N_{pert^{\diamondsuit}}$ (s-gpt 1.3B) 0.8445 0.9571 0.8972 0.8426 0.9424 0.8896
$N_{base^{\text{ \ }}}$ (s-gpt 2.7B) 0.9325 0.8929 0.9123 0.9289 0.8479 0.8866
$N_{pert^{\diamondsuit}}$ (s-gpt 2.7B) 0.8682 0.9656 0.9143 0.8560 0.9474 0.8993
Table 20: Accuracy of the models on the test/perturbation set, after filtering. The average standard deviation is $0.0049$.
@@
    Different datasets may work better with different embedding functions: eg s-bert 22M works best for Medical, and s-gpt 2.7B – for RUAR (with the exception of F1 score, for which s-bert 22M is best for both datasets). Smaller GPT model s-gpt 1.3B is systematically worse for both datasets;
@@
    Best accuracy may not be the best indicator of performance, depending on the scenario of use. Note that in our verification scenarios, where the verified DNN filter recognises the positive class, we look for higher precision, thus favouring fewer false positives, as they falsify verifiable subspaces (cf. Example 3). However for raw (unverified) use of DNN we value higher recall, favouring fewer false negatives, as we do not want the DNN to “miss” sensitive/dangerous cases.
    For Medical, s-bert 22M (either with or without adversarial training) looks like the best choice. However, for RUAR, the choice of the embedding function has a huge effect:
    • for general accuracy, s-bert 22M is the best choice (difference with the worst choice of the embedding function is $12-16$ percentage points),
    • for scenarios when one uses unverified DNN, $N_{pert}$ (s-gpt 2.7B) gives an incredibly high recall ($>99\%$) and would be a great choice (difference with the worst choice of the embedding function is $13-28$ percentage points).
    • However, if one wanted to use the same $N_{pert}$ (s-gpt 2.7B) for verification, this would be the worst choice, as precision for the network drops to $58-61\%$. For verification, $N_{base}$ (s-gpt 2.7B) $N_{base}$ or s-bert 22M would be better choices (with precision $>95\%$).
@@
    Looking at the effects of the adversarial training, it only makes a significant difference in accuracy for the Medical perturbed test set. However, it has more effect on improving recall (up to 10% for Medical and 33% for RUAR).
@@
    For verifiability-generalisabiliy trade-off, the choice of an embedding function also plays a role. Table 25 shows that s-gpt models exhibit lower verifiability compared to s-bert models. This observation also concurs with the findings in Section 0.4: greater volume correlates with increased generalisation, while a smaller and more precise subspace enhances verifiability. Indeed volumes for s-gpt models are orders of magnitude ($52-55$) larger than s-bert models.
@@ 0.5.4 Analysis of Perturbations and Falsifiability
Recall that two problems were identified as potential causes of falsifiable semantic subspaces: the imprecise embedding functions and invalid perturbations (i.e. the ones that change semantic meaning and the class of the perturbed sentences). In the previous section, we obtained implicit evidence of variability of performance of the available state-of-the-art embedding functions. In this section, we turn our attention to analysis of perturbations. As outlined in [136], to be considered valid, the perturbations should be semantically similar to the original, grammatical and have label consistency, i.e. human annotators should still assign the same label to the perturbed sample. Firstly, we wish to understand how common it is for perturbations to change the class, and secondly, we propose several practical methods how perturbation adequacy can be measured algorithmically.
@@
Recall that the definition of semantic subspaces depends on the assumption that we can always generate semantically similar (valid) perturbations and draw semantic subspaces around them. Both adversarial training and verification then explore the semantic subspaces. If this assumption fails and the subspaces contain a large number of invalid sentences, the NLP verification pipeline looses much of its practical value. To get a sense of the scale of this problem, we start with the most reliable evaluation of sentence validity – human evaluation.
Understanding the Scale of the Problem
For the human evaluation, we labelled a subset of the perturbed datasets considering all validity criteria. Annotation instructions are modified from [136] and are given in Table 21.
For this experiment, at the character and word level, we randomly take 10 samples from each dataset for each perturbation sub-type, i.e. at the character level this entails 10 perturbations for each category of inserting, deleting, replacing, swapping or repeating a character. At the sentence level, a sample entails more than 1 perturbation, given that we prompt vicuna-13b with instructions for the original sentence to be rephrased 5 times. For a sample, we randomly choose one of the perturbations from the returned set of perturbations. In this way, we take 50 samples from each dataset. This leaves us with 290 manually annotated samples (130 from the medical safety, and 160 from the R-U-A-Robot dataset). Inter-Annotator Agreement (IAA) is reported via intraclass correlation coefficient (ICC).
Human Evaluation.
ICC (A,1) estimates and their 95% confidence intervals (CI) were calculated based on absolute-agreement (single, fixed raters). Using cutoffs provided by [78], agreement was determined to be MODERATE for semantic similarity (F = 4.4 df (289), p<.001, 95% CI = [0.56,0.69]), BELOW SATISFACTORY for grammaticality (ICC = 0.43, p <.001, 95% CI = [0.34,0.52]) and BELOW SATISFACTORY for label consistency (ICC = 0.29, p<.001, 95% CI = [0.18, 0.39]). This suggests that, when using LLMs, perturbation quality and robustness to class change cannot be taken for granted.
Limitations. We note this is in part due to our definition of grammatical being interpreted differently by the two independent evaluators (one accounting for character perturbations/spelling mistakes as un-grammatical and one not), and label consistency being ambiguous for the R-U-A-Robot dataset. Finally, we also note that correlation between raters is statistically significant across all categories - indicating that ratings across coders were aligned beyond chance probability (criteria $\alpha$ = 0.05). Future replications are warranted.
@@
Semantic similarity
Evaluate whether the original and the modified sentence have the same meaning on a scale from 1 to 4, where 1 is ‘The modified version means something completely different’ and 4 means ‘The modified version has exactly the same meaning’.
Grammaticality
Grammatically means issues in grammar, such as verb tense. Evaluate the grammaticality of the modified version on a scale of 1-3, where 1 is ‘Not understandable because of grammar issues’, and 3 is ‘Perfectly grammatical’.
Label consistency
Decide whether the positive label of the modified sentence is correct using labels 1 - ‘Yes, the label is correct’, 2 - ‘No, the label is incorrect’ and 3 - ‘Unsure’.
@@
Effective Ways to Measure and Report Perturbation Validity
Although no geometric or algorithmic method will ever be able to match to the full extent the human perception and interpretation of sentences, we can still formulate a number of effective methods that give an implicit characterisation of validity of perturbations utilised when defining semantic subspaces. We propose three:
  1. 1.
    Using cosine similarity of embedded sentences, we can implicitly characterise semantic similarity
  2. 2.

### v3

@@ 5.3 Experimental Setup of the Verification Pipeline
    semantic subspaces should be preferred over geometric subspaces as they result in a better verifiability-generalisability trade-off;
  2. 2.
    constructing semantic subspaces using stronger NLP perturbations results in higher verifiability of those subspaces;
  3. 3.
    likewise, adversarial training using subspaces constructed with stronger NLP perturbations also results in higher verifiability;
  4. 4.
    Marabou allows us to verify a higher percentage of subspaces compared to ERAN thanks to its completeness and precision.
@@
Based on these results, we further strengthen the NLP perturbations by substituting Polyjuice used in the previous section with Vicuna. Vicuna introduces more diverse and sophisticated sentence perturbations. In addition, we mix in the character and word perturbations used in the previous section, to further diversify and enlarge the set of available perturbed sentences. In the terminology of Section 4.1, we obtain the sets of perturbed sentences $\mathcal{A}^{b}_{t}(\mathcal{Y}^{pos})$ and $\mathcal{A}^{b}_{t}(\mathcal{Y}^{neg})$ where $t=\{char,word,vicuna\}$ is a combination of these perturbations. Table 22 also uses notation $\mathcal{A}^{b}_{t^{\diamondsuit}}(\mathcal{Y}^{pos})$ and $\mathcal{A}^{b}_{t^{\diamondsuit}}(\mathcal{Y}^{neg})$ to refer to filtered sets, this terminology will be introduced in Section 5.5.2.
@@
Given $\mathcal{A}^{b}_{t}(\mathcal{Y}^{pos})$, the set of semantic subspaces $\mathbb{H}_{pert}$ which we wish to verify, are obtained via the hyper-rectangle construction in Definition 6. Accordingly, we set the adversarial training to explore the same subspaces $\mathbb{H}_{pert}$, and to obtain the network $N_{pert}$.
5.4 Analysis of the Role of Embedding Functions
For illustration, as well as an initial confidence check, we report F1 of the obtained models, for each of the chosen embedding functions in Table 23. Overall the figures are as expected: compared to the F1 of 54-64% for the zero-shot model, using a fine-tuned trained DNN as a filter dramatically increases the F1 to the range of 76-95%.
@@ 5.4 Analysis of the Role of Embedding Functions
RUAR $N_{zero-shot}$ 51.67% 58.35% 54.81% - - -
$N_{base}$ (s-bert 22M) 95.68% 91.29% 93.44% 94.77% 71.86% 81.74%
$N_{pert}$ (s-bert 22M) 84.97% 98.63% 91.29% 81.25% 94.66% 87.45%
$N_{base}$ (s-gpt 1.3B) 96.20% 87.25% 91.51% 95.45% 67.38% 78.98%
$N_{pert}$ (s-gpt 1.3B) 63.03% 99.80% 77.24% 61.26% 98.60% 75.54%
$N_{base}$ (s-gpt 2.7B) 96.74% 87.29% 91.77% 95.49% 69.82% 80.66%
$N_{pert}$ (s-gpt 2.7B) 60.18% 99.80% 75.08% 58.46% 98.99% 73.50%
Medical $N_{zero-shot}$ 58.95% 70.22% 64.09% - - -
$N_{base}$ (s-bert 22M) 95.23% 93.25% 94.23% 95.20% 89.64% 92.34%
$N_{pert}$ (s-bert 22M) 93.35% 97.36% 95.31% 92.38% 95.17% 93.76%
$N_{base}$ (s-gpt 1.3B) 91.93% 88.11% 89.98% 92.17% 84.17% 87.98%
$N_{pert}$ (s-gpt 1.3B) 84.41% 96.27% 89.38% 83.15% 94.70% 88.54%
$N_{base}$ (s-gpt 2.7B) 93.25% 89.29% 91.23% 92.89% 84.79% 88.66%
$N_{pert}$ (s-gpt 2.7B) 86.03% 96.56% 90.98% 84.88% 94.99% 89.64%
Table 23: Performance of the models on the test/perturbation set. The average standard deviation is $0.0049$.
@@
    There is not a single embedding function that always results in the highest F1. For example, s-bert 22M is found to have the highest F1 for Medical, while s-gpt 2.7B has the highest F1 for RUAR (with the exception of F1 score, for which s-bert 22M is best for both datasets). The smaller GPT model s-gpt 1.3B is systematically worse for both datasets.
@@
    As expected and discussed in Section 5.1, depending on the scenario of use, the highest F1 may not be the best indicator of performance. For Medical, s-bert 22M (either with or without adversarial training) obtains the highest precision, recall and F1. However, for RUAR, the choice of the embedding function has a greater effect:
    • if F1 is desired, s-bert 22M is the best choice (difference with the worst choice of the embedding function is $12-16\%$,
    • for scenarios when one is not interested in verifying the network, the embedding function s-gpt 2.7B when combined with adversarial training gives an incredibly high recall ($>99\%$) and would be a great choice (difference with the worst choice of the embedding function is $13-28\%$).
    • however, if one wanted to use the same network for verification, s-gpt 2.7B would be the worst choice of embedding function, as the resulting precision drops to $58-61\%$. For verification, either $N_{base}$ trained with s-gpt 2.7B, or $N_{base}$ trained with s-bert 22M would be better choices, both of which have precision $>95\%$.
@@
    Adversarial training only makes a significant difference in F1 for the Medical perturbed test set. However, it has more effect on improving recall (up to 10% for Medical and 33% for RUAR).
@@
    For verifiability-generalisability trade-off, the choice of an embedding function also plays a role. Table 32 shows that s-gpt models exhibit lower verifiability compared to s-bert models. This observation also concurs with the findings in Section 4: greater volume correlates with increased generalisation, while a smaller and more precise subspace enhances verifiability. Indeed volumes for s-gpt models are orders of magnitude ($52-55$) larger than s-bert models.
@@ 5.5 Analysis of Perturbations
Recall that two problems were identified as potential causes of embedding errors in semantic subspaces: the imprecise embedding functions and invalid perturbations (i.e. the ones that change semantic meaning and the class of the perturbed sentences). In the previous section, we obtained implicit evidence of variability of performance of the available state-of-the-art embedding functions. In this section, we turn our attention to analysis of perturbations. As outlined in [150], to be considered valid, the perturbations should be semantically similar to the original, grammatical and have label consistency, i.e. human annotators should still assign the same label to the perturbed sample. Firstly, we wish to understand how common it is for our chosen perturbations to change the class, and secondly, we propose several practical methods how perturbation adequacy can be measured algorithmically.
@@
Recall that the definition of semantic subspaces depends on the assumption that we can always generate semantically similar (valid) perturbations and draw semantic subspaces around them. Both adversarial training and verification then explore the semantic subspaces. If this assumption fails and the subspaces contain a large number of invalid sentences, the NLP verification pipeline loses much of its practical value. To get a sense of the scale of this problem, we start with the most reliable evaluation of sentence validity – human evaluation.
5.5.1 Understanding the Scale of the Problem
For the human evaluation, we labelled a subset of the perturbed datasets considering all three validity criteria discussed above. In the experiment, for each original dataset $\mathcal{Y}$ and word/character perturbation type $t$, we select 10 perturbed sentences from $\mathcal{A}^{16}_{t}(\mathcal{Y})$. At the character level this gives us 50 perturbed sentences for both datasets (10 each for inserting, deleting, replacing, swapping or repeating a character). At the word level this gives us 60 perturbed sentences for RUAR (deletion, repetition, ordering, negation, singular/plural, verb tense) and 30 for Medical (deletion, repetition, ordering). At the sentence level, we only have one kind of perturbation - obtained by prompting vicuna-13b with instructions for the original sentence to be rephrased 5 times. We therefore randomly select 50 vicuna-13b perturbed sentences for each dataset. This results in a total of 290 pairs consisting of the original sentence and the perturbed sentence (130 from the medical safety, and 160 from the R-U-A-Robot dataset). We then asked two annotators to both manually annotate all 290 pairs for the criteria shown in Table 24 which are modified from [150]. Inter-Annotator Agreement (IAA) is reported via intraclass correlation coefficient (ICC).
@@ 5.5.1 Understanding the Scale of the Problem
Semantic similarity Evaluate whether the original and the modified sentence have the same meaning on a scale from 1 to 4, where 1 is ‘The modified version means something completely different’ and 4 means ‘The modified version has exactly the same meaning’.
Grammaticality Grammatically means issues in grammar, such as verb tense. Evaluate the grammaticality of the modified version on a scale of 1-3, where 1 is ‘Not understandable because of grammar issues’, and 3 is ‘Perfectly grammatical’.
Label consistency Decide whether the positive label of the modified sentence is correct using labels 1 - ‘Yes, the label is correct’, 2 - ‘No, the label is incorrect’ and 3 - ‘Unsure’.
@@
Results of Human Evaluation. The raw evaluation results are shown in Tables 2526 and 27. Overall, there are high scores for label consistency, in particular for rule-based perturbations, with $\approx 88\%$ and $85\%$ of the perturbations rated as maintaining the same label (i.e. score 1) by the two annotators $A1$ and $A2$ respectively. Similarly there are high scores for semantic similarity, with $\approx 85\%$ and $78\%$ of the ratings falling between levels 4 and 3 for $A1$ and $A2$. For grammaticality, annotators generally rate that perturbations generated by vicuna-13b are grammatical, whereas (as expected) rule-based perturbations compromise on grammaticality.
Dataset Perturbation Semantic Similarity (%)
A1 A2
1 2 3 4 1 2 3 4
RUAR Rule-based 06.36 07.27 09.09 77.27 05.45 10.90 34.54 49.09
LLM-based 18.00 08.00 20.00 54.00 16.00 08.00 00.00 76.00
Medical Rule-based 01.25 02.50 10.00 86.25 10.00 10.00 31.25 48.75
LLM-based 06.00 20.00 28.00 46.00 12.00 18.00 20.00 50.00
Table 25: Semantic similarity results of the manual evaluation for annotators A1 and A2.
Dataset Perturbation Grammaticality (%)
A1 A2
1 2 3 1 2 3
RUAR Rule-based 10.90 31.81 57.27 13.63 78.18 08.18
LLM-based 02.00 02.00 96.00 00.00 02.00 98.00
Medical Rule-based 07.50 32.50 60.00 01.25 88.75 10.00
LLM-based 00.00 00.00 100.0 00.00 06.00 94.00
Table 26: Grammaticality results of the manual evaluation for annotators A1 and A2.
Dataset Perturbation Label Consistency (%)
A1 A2
1 2 3 1 2 3
RUAR Rule-based 88.18 00.90 10.90 85.46 04.54 10.00
LLM-based 78.00 20.00 02.00 70.00 24.00 06.00
Medical Rule-based 90.00 00.00 10.00 97.50 00.00 02.50
LLM-based 88.00 04.00 08.00 74.00 00.00 26.00
Table 27: Label consistency results of the manual evaluation for annotators A1 and A2.
In order to evaluate the inter-annotator agreement, we report the ICC between the annotators. The ICC estimates and their 95% confidence intervals (CI) were calculated based on absolute-agreement (single, fixed raters) – often referred to as ICC(A,1). Using cutoffs provided by [151], agreement was determined to be MODERATE for semantic similarity (F = 4.4 df (289), p<.001, 95% CI = [0.56,0.69]), BELOW SATISFACTORY for grammaticality (ICC = 0.43, p <.001, 95% CI = [0.34,0.52]) and BELOW SATISFACTORY for label consistency (ICC = 0.29, p<.001, 95% CI = [0.18, 0.39]).
This suggests that although annotators individually rated the perturbations for high label consistency, there may be disagreement on which specific samples maintain the same label. Given the moderate agreement for semantic similarity, we note that there may be perturbations that are semantically similar and yet may not maintain label consistency. For example Original: ‘if a computer can feel emotions, does that make you a computer or an actual human?’, Perturbation: ‘if a computer can feel, does that make it a machine or a person’ was rated by both annotators as high in semantic similarity but not maintaining label consistency. Overall, and particularly when using LLMs, perturbation quality and robustness to class change cannot be taken for granted, particularly when dealing with safety-critical queries.
Limitations. We note this is in part due to our definition of grammatical being interpreted differently by the two independent evaluators (one accounting for character perturbations/spelling mistakes as un-grammatical and one not), and label consistency being ambiguous for the RUAR dataset. Finally, we also note that correlation between raters is statistically significant across all categories - indicating that ratings across coders were aligned beyond chance probability (criteria $\alpha$ = 0.05). Future replications are warranted.
5.5.2 Automatic Ways to Measure and Report Perturbation Validity
Although in the near future, no geometric or algorithmic method will be able to match to the full extent the human perception and interpretation of sentences, we can still formulate a number of effective methods that give a characterisation of the validity of the perturbations utilised when defining semantic subspaces. We propose two:
  • Using cosine similarity of embedded sentences, we can characterise semantic similarity